### v1

# Error Estimation for Adaptive Mesh Refinement in Droplet Simulations

Abstract: We present a one-dimensional shear force driven droplet formation model with a flux-based error estimation. The presented model is derived using asymptotic expansion and a front-tracking method to simulate the droplet interface. The model is then discretized using the Galerkin finite element method in the mixed form. However, the jumps in the solution gradients are discontinuous and can grow faster due to the highly convective pinch-off process. This leads to an erroneous droplet interface and incorrect curvature. Therefore, the mesh must be sufficiently refined to capture the interface accurately. The mixed form of the governing equation naturally provides smooth interface gradients that can be used to compute the error estimate. The computed error estimate is then used to drive the adaptive mesh refinement algorithm.
1 Introduction
The formation of droplets and bubbles is a mature field in physics. These surface tension driven phenomena have been deeply explored using interface-tracking and interface-capturing approaches. Historically, the accurate description of free surface flows emerged after Laplace [1] and Young [2] explained the role of mean curvature and the surface tension force as a source of instability in 1805. Eggers and Villermaux presented a comprehensive review on understanding the liquid break-up process [3]. Starting with the initial fluid column, the drop becomes heavier by continuously adding fluid. Under an external force like gravity, the instabilities start to grow on the interface. Meanwhile, the surface tension minimizes the surface energy by decreasing the radius of the fluid column. A spherical droplet starts to form at the end of the fluid column hanging from a very thin neck. Eventually, the radius becomes zero and a droplet separates from the initial fluid column. The point of zero radius is called the “pinch-off" point. After the primary pinch-off, the neck recoils due to the “unbalanced” surface tension. The capillary waves due to this recoil perturb the interface before the tip can collapse back to the top of the fluid cone. This leads to a secondary breakup creating smaller droplets, which are called “satellite drops”.
Inaugural work on the mathematical modeling of droplet formation was done by Eggers and Dupont [4]. Their work involved finite difference formulation of the one-dimensional asymptotic model. Ambravaneswaran et al. [5] explored a finite element version of the same model. This model was only explored with the gravitational force as an external force. However, many industrial applications like atomization and droplet entrainment in annular flow [6, 7], and microfluidics devices [8, 9, 10] involves of droplet formation under a shear force environment. Using the same asymptotic approach for shear-induced droplets, we considered a one-dimensional model that estimates droplet sizes for an environment with co-flowing fluids [11]. For applications with a symmetrical structure of droplet formation, this one-dimensional model with the interface tracking approach is advantageous in terms of computational costs, while still maintaining the accuracy required for comparison with experiments. Nonetheless, this shear-induced droplet model deals with a wide range of external force magnitudes that induces a wide range of interface instability time frame.
The instability leading to the pinch-off point requires a finer mesh in the region of high curvature. As the interface evolves with the progression of numerical simulation, the mesh must be adaptively refined to control discretization errors, especially in the singularity regions. Therefore, a robust error estimation method is essential in designing an adaptive mesh refinement algorithm for our one-dimensional model. The field of a posteriori error estimation has been under continuous and rigorous improvement since the 1980s [12]. The preliminary work on error estimation was based on the finite element residuals and the flux jumps across the element boundaries [13, 14, 15, 16]. Since this work, there are many thorough review articles and books written on the topic of error estimation in finite element analysis from which we wish to refer here [17, 18, 19]. A thorough review on the recent advancements in the field of error estimation is in [20]. In the present work, we show that the mixed finite element formulation of our one-dimensional model naturally provides an effective error estimate for adaptive mesh refinement.
In the following sections, we describe the one-dimensional droplet pinch-off model and explain the need for an error estimator. Then, we explore the flux-based error estimation approach and show that the mixed form naturally provides a smooth flux to compute error bounds. Finally, we illustrate the usefulness of our error estimate using adaptive mesh refinement.
2 Problem Description
We consider a fluid column slowly flowing with the constant velocity $u^{d}$ out of a nozzle with the radius $h_{in}$. The continuous (or outer) fluid is co-flowing with the dispersed (or inner) fluid column but with a higher velocity $u^{c}$ in a capillary tube with radius $R$. Fig. (1) shows the schematic describing this co-flowing fluids scenario. The superscripts $d$ and $c$ represent dispersed phase and continuous phase fluids, respectively.
Figure 1: Schematic of a droplet formation in a co-flowing fluid.
We start with the Navier-Stokes equations in the cylindrical coordinates. The pinch-off process happens in finite time due to the surface tension forces trying to minimize the surface energy by contracting the interface in the radial direction. The radial contraction is faster than the axial elongation, which allows us to expand the solution variables asymptotically in $r$. Finally, the normal forces are balanced by the surface tension forces, which are used to simplify the equation into a one-dimensional model.
Similarly, we consider the droplet pinch-off in another co-flowing fluid under the gravitational and shear forces. We consider the co-flowing fluids with a preserved symmetry around the angular direction. This provides the basis for the one-dimensional approach. The continuous fluid affects the droplet morphology by applying shear force on the interface between two fluids. This shear force effect can be approximated with the asymptotically derived force balance on the interface, leading to the derivation of the one-dimensional governing equations [11]. The full derivation can be found in [21].
The momentum and the interface equations are given as follows.
$\frac{\partial u^{d}}{\partial t}+u^{d}\frac{\partial u^{d}}{\partial z}+\frac{\gamma}{\rho}\frac{\partial\mathcal{K}}{\partial z}-\frac{6\nu^{d}}{h}\frac{\partial u^{d}}{\partial z}\frac{\partial h}{\partial z}\left(1+\frac{\mu^{c}}{\mu^{d}}\right)-3\nu^{d}\frac{\partial^{2}u^{d}}{\partial z^{2}}\left(1+\frac{2}{3}\frac{\mu^{c}}{\mu^{d}}\right)$
$+\frac{2}{\rho^{d}}\frac{dp^{c}}{dz}+\frac{1}{2\rho^{d}\ln(C)}\frac{dp^{c}}{dz}-\left(1-\frac{\rho^{c}}{\rho^{d}}\right)g$ $=0$ (1)
$\frac{\partial h}{\partial t}+u^{d}\frac{\partial h}{\partial z}+\frac{h}{2}\frac{\partial u^{d}}{\partial z}$ $=0$ (2)
Here, parameters $\gamma$, $\rho$, $\nu$, $\mu$, and $p$ represent surface tension coefficient, density, kinematic viscosity, dynamic viscosity, and pressure, respectively. The superscript represents dispersed or continuous phase fluid. $\mathcal{K}$ is the curvature defined by
$\mathcal{K}=\left[\frac{1}{h\left(1+\frac{\partial h}{\partial z}^{2}\right)^{1/2}}-\frac{\frac{\partial^{2}h}{\partial z^{2}}}{\left(1+\frac{\partial h}{\partial z}^{2}\right)^{3/2}}\right]$ (3)
The quantity (C1)h(C-1)h defines the thickness of the shear layer in the continuous phase flow, which defines how much force is experienced by the dispersed phase droplet. The parameter $C$ is a free parameter that can be defined using curve-fitted monotonic functions on the numerical and experimental data. In [11], it was shown that this single parameter model can account for behavior in a wide range of dimensionless parameter values.
The system of governing equations given by Eqs. (1-2) are then discretized using a Galerkin finite element method using a mixed method [11]. The mixed finite element weak form is given by
$\int_{\Omega}q\left[\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial z}-\frac{6\nu}{h}\frac{\partial h}{\partial z}\frac{\partial u}{\partial z}\left(1+\frac{\mu^{c}}{\mu^{d}}\right)+\frac{\gamma}{\rho}\left\{-\frac{s\frac{\partial s}{\partial z}}{h\left(1+s^{2}\right)^{3/2}}-\frac{s}{h^{2}\left(1+s^{2}\right)^{1/2}}\right\}\right$.
$\qquad\left.+\frac{1}{2\rho^{d}\ln(C)}\frac{dp^{c}}{dz}-\left(1-\frac{\rho^{c}}{\rho^{d}}\right)g\right]d\Omega+\int_{\Omega}\nabla q\left[3\nu\left(1+\frac{2}{3}\frac{\mu^{c}}{\mu^{d}}\right)\frac{\partial u}{\partial z}\right$.
$\qquad\left.+\frac{\gamma}{\rho}\frac{\frac{\partial s}{\partial z}}{\left(1+s^{2}\right)^{3/2}}\right]d\Omega-\int_{\Gamma}q\left[3\nu\left(1+\frac{2}{3}\frac{\mu^{c}}{\mu^{d}}\right)\frac{\partial u}{\partial z}+\frac{\gamma}{\rho}\frac{\frac{\partial s}{\partial z}}{\left(1+s^{2}\right)^{3/2}}\right]d\Gamma=0$ (4)
$\int_{\Omega}v\left[\frac{\partial h}{\partial t}+u\frac{\partial h}{\partial z}+\frac{h}{2}\frac{\partial u}{\partial z}\right]d\Omega=0$ (5)
$\int_{\Omega}w\left[s-\frac{\partial h}{\partial z}\right]d\Omega=0$ (6)
where $s$ is the introduced variable for the mixed form that approximates the derivative of the interface distance, that is $\partial h/\partial z$.
The Eqs. (4-6) are solved using continuous Galerkin method with $C^{0}(\Omega)$ elements. The solution algorithm involves moving mesh and calculating the droplet length in a self-consistent way. The algorithm is explained in [22], and has been implemented using the PETSc libraries [23, 24].
The flow becomes highly convective as the droplet interface evolves up to a point when it starts forming the neck. The radial contraction of the interface accelerates faster than the elongation, hence rapidly approaching the singularity. This behavior should be captured accurately to get the precise equilibrium droplet profiles. In addition, the interface is advected with the flow, so that errors in the interface location at one step strongly affect later steps in the flow.
3 Error estimation
When approaching the singularity, the high-curvature regions quickly develop high stresses. An error in the solution can cause the interface to advect in an incorrect direction, providing an incorrect droplet profile. A coarser mesh must be refined to improve the solution in those regions and get an accurate force balance, leading to more accurate droplet profiles. This refinement can be done adaptively by estimating the true error in the solution using an a posteriori error estimation approach. The foremost goal is to compute the discretization error in the droplet interface evolution.
Since the solution is approximated with $C^{0}(\Omega)$ continuity, the gradients of the solutions are discontinuous across the element boundaries. This discontinuity strongly reflects on the curvature, which can produce erroneous equilibrium droplet profiles. We can use a flux-recovery based approach to drive the adaptive mesh refinement. The flux-recovery approach aims to post-process a smooth gradient from the FE solution and determine the error estimator based on the error in the smoothed and non-smoothed gradients.
In the droplet pinch-off model, we get the velocity ($u$) and the droplet radius ($h$) as the solution. However, the mixed form approach utilizes the approximation of the gradient $\partial h/\partial z$ as a part of the solution. In other words, the mixed variable $s$ already provides a smooth gradient of $h$. If the true gradient is $\partial\bar{h}/\partial z$, the quantity $s$ is assumed to retain better accuracy than $\partial h/\partial z$ since the slope (or flux) $s$ is part of the equilibrium equations and coupled with the curvature derivatives in the momentum equations. The true error norm in the gradient can be written as
[…]
The quality of the estimate clearly depends on how good the approximation ‘$s$’ is of the true slope $\partial\bar{h}/\partial z$. Assuming there exists $c\in(0,1)$ such that
[…]
Using the triangle inequalities, the bounds on the true error in terms of the error estimate follow directly that
[…]
The traditional a posterior error estimation methods like residual-based and flux-based methods use residual in each element, and flux jumps across the elemental interfaces. However, the error bounds achieved in these methods consist of an unknown multiplicative constant.
This approach can also be viewed from the residual-based error estimation perspective. A residual-based a posteriori error estimation approach directly utilizes the finite element solution. Defining the error in the energy norm, one can also derive upper and lower error bounds for the error in the slope ‘$s$’ using the element residuals of the Eq. (6).
Here, the error bounds given in Eq. (9) are used to refine the mesh adaptively. A simple $h$-refinement can be done cyclically:
solve $\rightarrow$ estimate error $\rightarrow$ mark elements $\rightarrow$ refine
We use two strategies to refine the mesh and compare their effectiveness.
(1) The elements are marked using the maximum threshold $\bar{\eta}=\lambda\ \max_{K\in\mathcal{T}_{h}}\eta_{K}$, where $\lambda\in[0,1]$. Here, $K$ represents an element from the triangulation $\mathcal{T}_{h}$. This simple strategy allows marking elements $K$ such that $\eta_{K}\geq\bar{\eta}$.
(2) The elements are marked using Dörfler’s method [25], where the elements are marked such that the total error in the marked elements is $\lambda$ times the total error in the entire domain.
4 Results
The droplet simulations presented in this section are for an 85% glycerol solution with an inlet velocity of 5 mm/s in the background with co-flowing air with 1 m/s velocity. The inlet radius of a nozzle is considered to be 2.5 mm and the radius of the outer capillary tube is 25 mm. The material properties are used from [22].
Figure 2: Change in the mean curvature with the droplet evolution.
The error estimation approach is entirely motivated to capture the regions with sharp changes in the mean curvature. Therefore, it is vital to understand the change in the mean curvature to set our expectations for error estimates. Fig. (2) shows a sequence of images showing the change in the mean curvature as the droplet evolves up to the primary pinch-off. Initially, the mean curvature is the largest at the droplet tip. The mean curvature starts increasing in the neck region of the fluid column as the droplet evolves. This is naturally due to the radial shrinkage of the column. Finally, the mean curvature is highest at the pinch-off location when the primary droplet separates from the fluid column. The regions with sharp curvature changes are the top of the neck, the region near the pinch-off, and the droplet tip. Therefore, a higher error measurement is expected in these regions. These images are from the simulation with adaptive mesh refinement for better illustration.
We first report the element-wise errors without performing adaptive mesh refinement. The droplet length and radius are non-dimensionalized using the inlet radius $h_{in}$ in all of the plots. The element-wise error estimate, denoted by $\eta_{K}$, is shown in Fig. (3). The error estimate is plotted for the initial droplet profile and an evolved droplet profile as shown in Fig. (3(a)) and Fig. (3(b)), respectively. The initial droplet profile has the highest curvature gradient closer to the tip. Thus, the highest error is detected at the tip, as the plot reports.
On the other hand, the evolved droplet has a large error at the top of the neck region. This error is associated with the incorrect interface in that region caught in the droplet profile. This inaccuracy leads to a divergence of residuals and simulation failure before reaching the primary pinch-off. The magnified image shows the error magnitude at the tip, which is larger for the last few elements. This suggests that mesh refinement is required to accurately represent a smooth spherical shape of the droplet at the tip.
(a) Initial droplet phase.
Refer to captionRefer to caption
(b) Evolved droplet phase with a magnified view at the tip.
Figure 3: Error plots showing the error magnitude at the initial phase and the evolved phase.
(a) Non-refined grid.
(b) Adaptively refined grid using Max strategy.
(c) Adaptively refined grid using Dörfler strategy.
Figure 4: Comparison of error plots superposed on the 85% glycerol droplet.
Based on the reported errors, the adaptive mesh refinement strategy as explained in the previous section is used for the simulations. As the simulation progresses, we track the error evolution along with the droplet interface. This provides an insight into how the error progresses as the curvature changes. Using the maximum strategy and the Dörfler’s strategy, adaptive mesh refinement is performed when the droplet length is greater than $N$ times the inlet radius, where $N=\{2,3,4,$...\}. The mesh is adaptively refined with $\lambda=0.1$. This means that the elements where the error is larger than 10% of the maximum value are marked and refined. For the Dörfler’s approach, the elements with a cumulative error equal to 90% of the total error are marked and refined. The choice of these criteria is based on the observation in the error evolution for a non-refined grid.
Fig. (4) shows the evolved error superposed on the droplet profile for both no refinement and adaptive refinement. The evolution profiles are shown at the same time stamp. Fig. (4(a)) shows the error plot along the droplet length on a non-refined grid. The top of the neck is a high-velocity region where the upstream flow pushes the droplet axially. However, the radial velocity is much faster than this axial push. Hence as explained before, the discontinuous interface slope (or the gradient) grows quickly. This is visible at the top of the neck where the element-wise error is large. Due to the large error, the droplet interface shows instabilities and hence the simulation diverges. This instability is visible in the droplet profile.
Fig. (4) shows the comparison of the error on non-refined mesh, refined mesh with maximum strategy, and refined mesh with Dörfler’s approach. Compared to the non-refined grid, the error on a refined mesh shows an error reduction in the order of millions. The curvature at the top of the neck region is much smoother, and the interface instabilities are absent. Comparing Fig. (4(b)) and Fig. (4(c)), the Dörfler’s approach requires fewer elements to refine compared to the maximum strategy. The resulting element-wise error in the Dörfler’s strategy case is still on the order of $10^{-6}$.
While comparing only the droplet profile, the non-refined mesh is coarse at the tip, whereas the refined version of the droplet is much smoother. The droplet profiles in mesh with the maximum strategy and the Dörfler’s strategy are almost identical. It is crucial to have an accurate droplet profile to calculate the volume of the pinched-off droplet. The refined version provides accuracy and stability with optimally reduced computational cost.
5 Conclusion
We presented a simple yet effective strategy to compute the error estimate for droplet simulations with front tracking method. The one-dimensional model used for the simulations is accurate for shear-induced droplets in a co-flowing environment. The mixed form equations naturally incorporate a smooth flux as a part of the solution, which allows us to use the flux-based error estimation. The comparison of error evolution for 85% glycerol droplets in quiescent air is presented for a non-refined and adaptively refined grid. This comparison shows that the adaptively refined grid provides improved stability and accuracy in capturing the droplet interface. The one-dimensional model has been previously used for conducting uncertainty quantification (UQ) analysis for droplet atomization in hybrid rocket combustion [26]. The presented error estimate can further enhance the accuracy of the UQ analysis.